### v1

# Relevance for Human Robot Collaboration

Abstract
Effective human-robot collaboration (HRC) requires the robots to possess human-like intelligence. Inspired by the human’s cognitive ability to selectively process and filter elements in complex environments, this paper introduces a novel concept and scene-understanding approach termed ‘relevance.’ It identifies relevant components in a scene. To accurately and efficiently quantify relevance, we developed an event-based framework that selectively triggers relevance determination, along with a probabilistic methodology built on a structured scene representation. Simulation results demonstrate that the relevance framework and methodology accurately predict the relevance of a general HRC setup, achieving a precision of 0.99 and a recall of 0.94. Relevance can be broadly applied to several areas in HRC to improve task planning time by 79.56% compared with pure planning for a cereal task, reduce perception latency by up to 26.53% for an object detector, improve HRC safety by up to 13.50% and reduce the number of inquiries for HRC by 75.36%. A real-world demonstration showcases the relevance framework’s ability to intelligently assist humans in everyday tasks.
I Introduction
Robots and automated systems are increasingly integrated into modern society, and the human world, which requires them to develop human-like intelligence and cognitive abilities [1, 2]. One example of such abilities is that human beings have the capability and tendency to focus on the relevant physical elements in an environment and relevant abstract concepts or information in the input based on the context, the objectives, the historical experiences, the reasoning, etc [3]. This ability of humans is demonstrated in various studies [4, 5].
Empowering robots with this human-like cognitive capability offers two significant benefits. First, robots interacting closely with human beings will achieve a scene-understanding similar to that of humans. This results in more natural and effective behavior and interaction. Second, by identifying the most relevant elements in the scene, the robot can effectively focus its limited computation power on the relevant elements to reduce the computation requirements, increase the computation speed, and even increase the safety of the robots in HRC tasks [3].
Previous works have primarily focused on visual saliency and attention mechanisms to identify prominent features. However, they often overlook the broader context necessary for long-term task planning and proactive assistance. Additionally, attention mechanisms are typically constrained to short-term simple motion generation and lack interpretability, which makes them unsuitable for complex problems requiring comprehensive reasoning. These limitations highlight the need for a novel approach designed for context-aware, efficient, and accurate scene understanding, reasoning about the interconnection among humans, dynamic environments, context, etc.
In this paper, we define and introduce a novel concept and approach for scene understanding, termed ‘relevance’. In relevance, our approach to scene understanding starts by identifying all elements in the scene, which are then grouped into classes. Our approach identifies the classes of contextual elements pertinent to what is or may happen in the scene. Our approach helps identify all classes with supplementary elements. Such classes/elements are not core to the scene with a positive (enhancing), negative (distracting), or neutral (optional) effect. Finally, the approach identifies the core or essential elements, focusing on indispensable components like main characters, primary actions, and key elements for what is taking place. This approach ensures a structured, coherent, systematic, and efficient scene understanding that is updated in real time.
To quantify relevance, we propose and develop a novel framework. The event-based framework contains modules of perception, triggers, relevance determination, and decision-making. The event-based mechanism initializes the relevance determination only under certain circumstances, resulting in a dramatic increase in computation speed. In the relevance determination module, a unified probabilistic methodology based on a novel scene representation is proposed and developed. This methodology accurately quantifies the relevance of the scene, which dramatically improves the performance of HRC.
To summarize, the contributions of this paper are four-fold: 1. We proposed a new concept and scene understanding approach called relevance to identify the relevant elements in the scene for the human’s objective. 2. We developed an event-based framework for faster relevance determination. 3. We developed a systematic methodology for relevance quantification and determination. 4. We demonstrated the effectiveness of our framework and methodology using both simulation evaluation and real-world demonstration.
II Related Works
Two areas of research, visual saliency and attention, share similar goals to our work but are fundamentally different. To the best of our knowledge, this work is the first in HRC to empower robots with cognitive ability by identifying and focusing on the relevant elements based on long-term objectives and context.
II-A Visual Saliency
Visual saliency seeks to identify and localize the most distinctive regions or objects within an image based on visual conspicuity without considering the context. Traditional methods are developed based on the features at the pixel level, including colors, intensity, orientation, etc [6, 7, 8]. Deep learning-based methods can also be applied [9, 10, 11, 12]. However, visual saliency only considers visual conspicuity, neglecting important considerations, such as context, for proactive assistance in HRC.
II-B Attention
Attention is another mechanism that selectively focuses on relevant parts of the input data but is dramatically different from relevance. In a nutshell, attention is structured with self-attention and/or cross-attention to formulate and optimize weights for the input features [13]. However, attention in robotics is only applied to end-to-end algorithms for better completion of the current short-term action, e.g., pick and place [14, 15, 16]. Moreover, attention has several challenging limitations, such as limited applicability only to end-to-end methods, limited interpretability, and myopia. Relevance is a fundamental function in robot intelligence for high-level human-like reasoning by considering human factors (long-term objectives, goals, preference, and motion), task models (task structure, task prediction, object applicability), environmental factors, etc.
From another perspective, applying attention to human-robot collaboration (HRC) is challenging due to three factors: the proactive nature of HRC, its inherent complexity, and the difficulty of dataset collection. First, attention mechanisms are effective in reactive tasks where the neural networks focus on specific parts of the immediate input to generate actions. However, for proactive assistance in general HRC, the model is supposed to consider long-term goals, which require projection into the future. Second, the complexity of HRC is very high, where the robots are supposed to reason the interconnection between the context, the tasks in a hierarchy, and the human (objective and preferences), which is challenging for attention mechanisms to capture. Third, end-to-end learning, where attention works, requires intensive data collection, which is challenging because of the dynamics and the higher dimension of the environment involving human actions.
III Framework for Relevance
In this section, we introduce our novel event-based framework for relevance determination and robotic sensorimotor action generation, as shown in Fig. 1. The framework contains perception module, triggers $T$ check, relevance determination, and decision making.
Figure 1: Overview of the framework for relevance quantification and application of relevance for proactive human-robot collaboration. The framework consists of a continuously running perception module, a trigger check module to selectively initialize relevance determination, a two-level relevance determination methodology, and a decision-making module to generate natural and efficient human robot interaction.
III-A Perception module
The perception module continuously processes the sensor input (camera, microphone, etc.) and obtains the features $\mathcal{F}$ of the scene. Unlike other pipeline frameworks for robots that process the perception and decision-making in a serial manner [17], our developed and proposed framework perceives the scene in an uninterrupted manner.
III-B Triggers Check
The triggers $T$ are to determine if a subprocess should be initialized for relevance determination based on the features $\mathcal{F}$. Some potential criteria for trigger activation include element changes and/or updated objectives or new tasks for the human.
III-C Relevance Determination
Our relevance determination methodology is a probability framework based on a new type of scene representation, i.e. classes and elements. The relevance determination methodology is a two-step methodology that determines the relevance classes and then the relevance elements. At the same time, the methodology determines if sufficient information is available. With sufficient information, the framework will go through the methodology and determine relevance of the scene. With information insufficient, the algorithms will request or wait for the perception module to give out more information and cues. More details of our methodology for relevance determination will be discussed in Section IV.
III-D Decision Making
The output of the relevance determination module will be utilized in the decision-making to provide natural, faster, and safer actions for proactive HRC.
IV Methodology for Relevance
In this section, we introduce the methodology for the relevance determination module in Fig. 1.
IV-A Class and Element Scene Representation
To quantify relevance efficiently and systematically, we propose a new representation $\mathcal{S}$. In $\mathcal{S}$, the scene is represented as a set of classes of elements:
$\mathcal{S}=\{C_{1},C_{2}\dots,C_{m}\}$ (1)
where $C_{i}$ represents the $i$th class of the classification results of all elements in $\mathcal{E}$. $\mathcal{E}$ is the set of all elements in the scene. $m$ is the number of classes.
Each $C_{i}$ contains a set of elements classified into the class $C_{i}$ according to the criterion $k_{i}\in k$, where $k=\{k_{1},k_{2},\dots,k_{m}\}$ is the set of classification criteria for each class. Thus, the set of elements belonging to the class $C_{i}$ based on criterion $k_{i}$ can be mathematically represented as:
$C_{i}=\{e_{j}\mid e_{j}\in\mathcal{E},e_{j}\text{ satisfies criterion }k_{i}\}.$
Each element $e_{j}$ is included in the class $C_{i}$ if it satisfies the criteria $k_{i}$.
In organizing elements, we consider two primary types of classes in this work: attribute classes and functional classes. Attribute classes group elements based on shared characteristics or properties, such as similar semantics or attributes. For example, attribute-based classes include general groupings like a class of fruits or books. Other examples within attribute classes are hierarchical classes, temporal classes, spatial classes, quantitative classes, etc. Functional classes, on the other hand, group elements based on their roles, actions, or contributions to a common objective. Examples include behavioral classes, conceptual classes, and thematic classes.
IV-B Relevance Determination Mechanism
We developed and proposed a multi-level relevance determination mechanism that works on the class level first and then on the element level within the relevant classes later. The relevance mechanism in both class and element level is defined as $\mathcal{M}$. The mechanism $\mathcal{M}$ takes the set of classes or elements and the set of features $\mathcal{F}$ as inputs. $\mathcal{F}$ can be represented as =$\mathcal{F}=\{\mathcal{F}_{v},\mathcal{F}_{a},\mathcal{F}_{c},\dots)$, where $\mathcal{F}_{v}$, $\mathcal{F}_{a}$, and $\mathcal{F}_{c}$ represent visual (e.g., elements detected, classes classified, motion information, etc.), auditory (e.g. conversation in the scene, etc.), and contextual (e.g., objective, tasks, etc.) features, respectively.
There are two types of outcomes from the mechanism $\mathcal{M}$. Suppose the information is sufficient for the relevance quantification and determination. In this case, the outcomes are the relevance scores for each class or element, and, at the same time, the set of relevant classes $\mathcal{C}_{r}$ or the set of relevant elements $\mathcal{E}_{r}$. If the information in $\mathcal{F}$ is not sufficient for relevance quantification and determination, then the algorithm will refer to the perception module to either actively request or obtain new cues $\mathcal{F}_{new}$ from the scene before continuing the processing. Thus, the mechanism $\mathcal{M}$ can be mathematically represented as:
$\mathcal{M}:\mathcal{F},\mathcal{X}\rightarrow\left\{\begin{array}[]{ll}\mathcal{R}(x)\text{ for }\forall x\in\mathcal{X}\text{, and }\mathcal{X}_{r}&\text{if }\text{sufficient }\mathcal{F}\\ \text{ask or wait for }\mathcal{F}_{new},&\text{else}\end{array}\right.$ (2)
where $\mathcal{X}$ represents a set of classes or elements, $x$ represents a specific class or element, and $\mathcal{X}_{r}$ represents a set of relevant classes or elements.
Both on the class level and the element level, the relevance $\mathcal{R}(x)$ is defined as the probability of $x$ being relevant given the objective of the human $\mathcal{O}$:
$\mathcal{R}(x)=P(x\mid\mathcal{O})$ (3)
(3) can also be conditioned on other factors according to the requirements of the applications. After $\mathcal{R}(x)$ is quantified, there are two possibilities: $\mathcal{R}(x)$ is larger than a threshold $\tau_{x}$ implies that $x$ is relevant in the scene to the objective, and then $x$ is added to the relevant set $\mathcal{X}_{r}$:
$R(x)>\tau_{\text{x}}\rightarrow x\text{ is relevant}\rightarrow\mathcal{X}_{r}=\mathcal{X}_{r}\cup x$ (4)
If $\mathcal{R}(x)$ is smaller than $\tau_{x}$, then $x$ is deemed as irrelevant:
$R(x)<\tau_{\text{x}}\rightarrow x\text{ is irrelevant}$ (5)
(3) can be influenced by a comprehensive set of possible factors. In this paper, we compute (3) based on four factors: the human’s objective, tasks, preferences, and spatial placement.
IV-C Class Relevance Modeling
In this part, we model the class relevance based on (3), which is defined as:
$\mathcal{R}(C_{i})=P(C_{i}\mid\mathcal{O})$ (6)
There could be several possible tasks $\mathcal{T}$ associated with the objective $\mathcal{O}$. Thus, we use the total probability to decompose (6) and obtain:
$\mathcal{R}(C_{i})=\sum_{\mathcal{T}}P(C_{i}\mid\mathcal{T},\mathcal{O})P(\mathcal{T}\mid\mathcal{O})$ (7)
In (7), the first term $P(\mathcal{C}_{i}\mid\mathcal{T},\mathcal{O})$ represents the probability that class $C_{i}$ is relevant given the task is $\mathcal{T}$, and the objective is $\mathcal{O}$. The second term $P(\mathcal{T}\mid\mathcal{O})$ represents the probability that the current task is $\mathcal{T}$ given the objective is $\mathcal{O}$. If the human’s historical preference data about the specific class $C_{i}$ for the task $\mathcal{T}$ and objective $\mathcal{O}$ is available, then the first term can be derived as the probability of $C_{i}$ is relevant based on the preference. If the historical preference data is not available for class $C_{i}$ for task $\mathcal{T}$, the probability $P(C_{i}\mid\mathcal{T},\mathcal{O})$ can be derived from predictions based on Large Language Models (LLM), other datasets that incorporate action sequences, other human cues, etc. The second term in (7) can be derived from prediction models for human tasks.
If any term in (7) is not available, or available, but with high uncertainty, the information to determine the relevant class is deemed as not sufficient, and $\mathcal{F}_{new}$ is required for accurate quantification of the class relevance. The uncertainty of the term $P(\mathcal{T}|\mathcal{O})$ can be estimated using entropy as:
$H(P(\mathcal{T}|\mathcal{O}))=-\sum_{\mathcal{T}}P(\mathcal{T}|\mathcal{O})\ln{P(\mathcal{T}|\mathcal{O})}$ (8)
where $H$ represents the entropy of a probability.
IV-D Element Relevance Modeling
With a similar methodology to class relevance modeling, the relevance for an element can be defined and modeled as:
$\mathcal{R}(e_{j})$ $=P(e_{j}|\mathcal{O})$ (9)
$=P(e_{j}|C_{i},\mathcal{O})P(C_{i}|\mathcal{O})$
$\phantom{=}+P(e_{j}|\neg C_{i},\mathcal{O})P(\neg C_{i}|\mathcal{O})$
where $e_{j}\in C_{i}\in C_{r}$, and the symbol $\neg$ denotes logical negation. $P(e_{j}|\neg C_{i},\mathcal{O})$ represents the probability that $e_{j}$ is relevant given the class $e_{j}$ belongs to is not relevant and the objective is $\mathcal{O}$, which equals to 0. Thus, (9) becomes:
$\mathcal{R}(e_{j})=P(e_{j}|C_{i},\mathcal{O})\mathcal{R}(C_{i})$ (10)
One unique consideration when modeling element relevance, but not for class relevance, is that some elements may share the same attributes except for the spatial factors. Thus, we model our element relevance based on a two-hierarchy approach in which each element is characterized by top-level attributes, such as type, and low-level attributes, such as spatial or temporal factors. With this consideration, (10) can be computed as:
$\mathcal{R}(e_{j})$ $=\mathcal{R}(C_{i})P(e_{j}|C_{i},\mathcal{O})$ (11)
$=\mathcal{R}(C_{i})(P(e_{j}|h_{j},C_{i},\mathcal{O})P(h_{j}|C_{i},\mathcal{O})$
$\phantom{=}+P(e_{j}|\neg h_{j},C_{i},\mathcal{O})P(\neg h_{j}|C_{i},\mathcal{O}))$
$=R(C_{i})P(e_{j}|h_{j},C_{i},\mathcal{O})P(h_{j}|C_{i},\mathcal{O})$
where $h_{j}$ represents the event that the high-level attributes of $e_{j}$ are relevant. The value $P(e_{j}|\neg h_{j},C_{i},\mathcal{O})=0$ in the above equation.
When determining if the information is sufficient for relevant element determination, the availability and/or uncertainty of the three terms in (11) should be carefully examined similarly to relevant class determination.
IV-E Constraints and Dependencies
Constraints and dependencies among elements, such as temporal and spatial constraints, can be considered in a post hoc manner for the downstream decision-making modules after relevance is determined. If a relevant element depends on or is constrained by an element, this element should also be considered in relevance and decision-making. It is critical to consider these constraints so that the downstream decision-making module can be informed comprehensively and produce better actions.
V Simulation Evaluation
A comprehensive simulation evaluation is conducted in this paper for the aspects of relevance quantification and its application to HRC.
V-A Simulation Domain Setup
We introduced two distinct HRC testing domains for our evaluations with detailed description as follows:
Coffee: The Coffee domain involves making coffee with creamer and passing it to humans. The original simple problem setup includes 19 objects with no task dependencies. In randomly generated cases, 10 to 30 elements from 200 irrelevant kitchenware items are added, along with three random spatial constraints introducing task dependencies. The hard problem setup starts with 36 elements, and randomly generated cases add 20 to 50 more, with eight spatial constraints. In both cases, the constraints require moving an object on top of another to access the lower one.
Cereal: The Cereal domain focuses on a robot preparing cereal, serving it to a person, and returning the items to their original places. In the original simple problem, there are 18 elements and 20 natural task dependencies, such as needing to open a cabinet to access the cereal. In random problem instances, 10 to 45 irrelevant kitchenware items are added. The hard version starts with 26 objects and 20 task dependencies, with an additional 30 to 60 irrelevant objects added in random cases without new task dependencies.
V-B Relevance Quantification Evaluation
For the relevance quantification evaluation, we adopted and developed several metrics. First, we define precision $\mathfrak{P}$, recall $\mathfrak{R}$, and $F_{1}$ score $\mathfrak{F}$, which evaluate the prediction of relevant element set based on the ground truth of the relevant element set. Moreover, we introduced a metric, $\mathfrak{N}$, representing the number ratio of predicted relevant elements to actual relevant elements for assessing speed improvement in downstream tasks. A higher $\mathfrak{N}$ will probably slow down the downstream tasks. Those metrics systematically evaluate the relevance quantification from aspects of effectiveness, completeness, and efficiency. Different applications will require different combinations of optimal metrics.
In this paper, we emphasize the simulation for the coffee domain for the brevity and clarity of our analysis. In the simulation, the objective is predefined as “get something to drink at the break of a conference”, and the task is defined as “drink cold brew coffee”. A large language model (LLM) is adopted to classify all the elements into the scene representation $\mathcal{S}$ and provide the probabilities necessary for the relevance quantification. We tested 25 combinations of the class thresholds $\tau_{c}$ and element thresholds $\tau_{e}$, with 30 simple cases randomly generated for each combination.
The simulation results are shown in Fig. 2, which agree well with common sense. As $\tau_{c}$ and $\tau_{e}$ increase, more classes and elements will be pruned. Thus, the recall decreases, the precision increases, and the relevant object ratio decreases as these two thresholds increase. the F1 scores demonstrate a trend that increases first and decreases with the threshold increase.
Figure 2: Simulation results for the coffee domain. The values in the figures are averaged across 30 cases for each threshold combination. When $\tau_{c}$ and $\tau_{e}$ equal to 0.2, our methodology archives a recall $\mathfrak{R}$ of 0.94, a precision $\mathfrak{P}$ of 0.99, an F1 score $\mathfrak{F}$ of 0.96, and an object ratio $\mathfrak{N}$ of 0.94.
The trend of $\mathfrak{R}$ in Fig. 2(a) demonstrates that when $\tau_{c}$ and $\tau_{e}$ equal to or are smaller than 0.2, most relevant elements can be retained. At the same range of the thresholds, according to Fig. 2(d), the number of relevant elements predicted is very close to the number of relevant elements in the ground truth, which is very small when compared with the total number of elements in the scene. This demonstrates that our relevance quantification methodology can successfully remove most irrelevant elements while preserving relevant elements.
Figure 3: Number of cases each element is predicted to be relevant out of 30 cases without considering constraints. $\tau_{c}$ and $\tau_{e}$ are 0.2. Our relevance quantification method predicts relevance accurately and reliably.
We select the thresholds $\tau_{c}$ and $\tau_{e}$ to be 0.2. The number of cases in which each type of element is relevant is shown in Fig. 3 without considering the constraints. Of the 30 cases, a couple randomly add an ice bucket to the setup and predict that the ice bucket is relevant. All 30 cases predict that coffee and plastic cups are relevant. These results further demonstrate that relevance prediction is very precise.
V-C Relevance in HRC
One benefit of relevance in HRC is that relevance can help the robot better reason about the scene and the human’s requirements in a well-structured manner, which resembles how humans interact with other humans. This benefit can be demonstrated with a simple robot inquiry example in a coffee serving setup. The human-like robot will first detect the relevant classes and types of elements. If the relevance of any classes or types of elements is high, those classes or types are deemed necessary (coffee in this domain). Other relevant classes or types of elements are optional, and the robot asks the human if they need those relevant elements. The number of inquiries required using different methods is shown in Table I. It is shown that with precise and proper reasoning about the relevance of an element to a task or objective, the number of inquiries required for appropriate assistance is reduced by 75.36% compared with the number without relevance. Thus, human-robot interactions can be much more natural and fluent.
TABLE I: Number of inquiries required for appropriate assistance. Relevance dramatically reduces the number of inquiries.
method inquiry elements count
relevance
with necessity
creamer, plastic cup, paper cup,
whole milk, reduced milk, stick, napkin
9
relevance
coffee, creamer, plastic cup, paper cup,
whole milk, reduced milk, stick, napkin
10
no relevance Everything in the scene 36.53
V-D Relevance in Task Planning
This section demonstrates the effectiveness of relevance in task planning. A Fast Downward planner is employed. We compare three methods: our relevance with 0.2 as thresholds, random relevance with 0.2 as thresholds, and pure planning without relevance. For each method, we evaluate the processing time (relevance computation + planning), the timeout rate (120s), and the failure rate. Results are shown in Table II.
Relevance dramatically improves task planning performance. Compared with Pure Planning, relevance significantly reduces the time cost by up to 79.56% and the timeout rate across all problem formulations. As the problem formulation becomes more complex, we observe a notable increase in the timeout rate and planning time for Pure Plan, highlighting its inefficiency in solving complex task-planning problems. When compared with random relevance, we observe a substantial reduction in failure rates, demonstrating that our relevance successfully and accurately determines the relevant elements of the scene.
TABLE II: Task Planning Performance Comparison. Our methodology of relevance determination accurately predicts relevance and dramatically reduces the planning time.
Domains Metrics Relevance
Pure
Planning
Random
Relevance
Coffee simple Time 16.76 45.84 21.61
Timeout 0.00 0.22 0.05
Fail 0.04 0.00 0.62
Coffee hard Time 36.91 - 21.38
Timeout 0.00 1.00 0.02
Fail 0.01 0.00 0.88
Cereal simple Time 13.37 63.84 39.61
Timeout 0.00 0.38 0.05
Fail 0.02 0.00 0.58
Cereal hard Time 30.70 - 53.82
Timeout 0.00 1.00 0.12
Fail 0.04 0.00 0.60
V-E Relevance in Perception
By selectively processing the regions containing relevant elements, a notable maximum reduction of 30.09% in inference time and 26.53% in total time per frame of an object detector is achieved. Additionally, relevance improves two safety metrics by 11.25% and 13.50%, respectively [3].
VI Experimental Demonstration
In this section, we present the real-world demonstration to verify the effectiveness of our proposed framework and relevance determination methodology.
VI-A Experimental Setup
The experimental setup is shown in Fig. 4(a). On one end of the table, a UR5 robot arm with a robotiq gripper is mounted to reason about the relevance and generate actions to assist humans. The table is a snack table for a conference with desserts, drinks, and utensils. A microphone is placed on the table to pick up the audio information and cues in the scene. The HRC task is for the robot to assist two humans into the scene one by one. The robot utilized relevance for optimal decision-making and HRC.
VI-B Demonstration Results
The demonstration results are shown in Fig. 4(b-h). In Fig. 4 (b), in the absence of a human in the scene, due to the lack of triggers, the perception module is continuously running, and no relevance determination is initialized. In Fig. 4(c-d), a human agent with recorded preferences for coffee enters the scene and grabs a coffee. The trigger is activated, and relevance is determined. Based on the recorded preferences, the robot actions are automatically generated to serve the human agent whole milk and a stir stick. In Fig. 4(e), a second human enters without indicative actions. The trigger is activated, but the relevance determination module lacks sufficient cues. The system continues looping through the perception module to collect more information. In Fig. 4(f-g), a conversation between two human agents about coffee activates the trigger to determine relevance. Without preference for the second human, the robot first generates actions related to the necessary elements (i.e., the coffee) and then inquires about the second human’s preferences for making the coffee. Based on the human response, the system generates a complete action sequence to serve the second human agent. In Fig. 4(h), two humans successfully get their preferred coffee with the robot’s assistance, demonstrating the effectiveness of relevance.
Figure 4: The illustration of (a) experimental setup and (b-h) demonstration results. With relevance, the robot fluently and successfully assists two human agents with cold-brew coffee drinking.
VII Conclusion
This paper presents a novel concept and scene-understanding approach, termed ’relevance,’ designed to identify pertinent elements within a scene to better support human objectives. To quantify relevance, we developed an innovative framework that continuously monitors the scene and selectively activates relevance assessment. We developed a probabilistic methodology with a novel scene representation to efficiently and accurately quantify relevance. Simulations demonstrate the framework’s effectiveness, achieving a precision of 0.99 and a recall of 0.94. to efficiently and accurately determine relevance. Simulation results demonstrate the effectiveness of this framework and methodology, achieving a precision of 0.99 and a recall of 0.94. Relevance can significantly enhance HRC by improving task planning time by 79.56% compared to pure planning for a cereal task, reducing perception latency by up to 26.53% for object detection, enhancing safety by up to 13.50%, and decreasing the number of inquiries by 75.36%. The experimental results show a significant advancement toward creating a versatile robotic assistant capable of providing intelligent support to humans.